OpenStudy (anonymous):

Two cars leave town going opposite directions. One car is traveling 55 mph, and the other is traveling 65 mph How long will it take before they are 180 miles apart? Hint: The time for both cars is the same and can be represented by "t." The total distance is 180 miles. The distance = (rate)(time). If you add the (rate)(time) of the first vehicle to the (rate)(time) of the second vehicle, that will equal the total distance of 180 miles. Since you only have one unknown (t), you only need one equation.

OpenStudy (anonymous):

Assuming they start at the same time and location, use the expression \[55t+65t=180\]

OpenStudy (anonymous):

55 + 65 =120 D=180 T=180/120=1.5 HRS.

OpenStudy (anonymous):

yes